# Adversary Resilient Learned Bloom Filters

### v1

@@ Abstract
Creating an adversary resilient Learned Bloom Filter [1] with provable guarantees is an open problem [2]. We define a strong adversarial model for the Learned Bloom Filter. We also construct two adversary resilient variants of the Learned Bloom Filter called the Uptown Bodega Filter and the Downtown Bodega Filter. Our adversarial model extends an existing adversarial model designed for the Classical (i.e not “Learned”) Bloom Filter by Naor Yogev [3] and considers computationally bounded adversaries that run in probabilistic polynomial time (PPT). We show that if pseudo-random permutations exist, then a secure Learned Bloom Filter may be constructed with $\lambda$ extra bits of memory and at most one extra pseudo-random permutation in the critical path. We further show that, if pseudo-random permutations exist, then a high utility Learned Bloom Filter may be constructed with $2\lambda$ extra bits of memory and at most one extra pseudo-random permutation in the critical path. Finally, we construct a hybrid adversarial model for the case where a fraction of the workload is chosen by an adversary. We show realistic scenarios where using the Downtown Bodega Filter gives better performance guarantees compared to alternative approaches in this hybrid model.
@@ 1 Introduction
The Bloom Filter is a probabilistic data structure that solves the Approximate Membership Query Problem. The data structure now known as the “Bloom” Filter was initially proposed as method $2$ in the section “Two Hash-Coding Methods with Allowable Errors” in a 1970 paper by Burton H. Bloom [4][5]. The Bloom Filter has applications in databases, cryptography, computer networking, social networking [6], and network security [7]. The Learned Bloom Filter is a novel data structure first proposed by Kraska et al [1] in 2017. The Learned Bloom Filter can be thought of as a Bloom Filter working in collaboration with a Learning Model. There are currently no known provably secure constructions of the Learned Bloom Filter. Prior work [2] has left the security of the Learned Bloom Filter as an open problem.
@@ 1.1.1 Classical Bloom Filter
A Bloom Filter for representing set $S$ with cardinality $n$ is a zero initialized array of $m$ bits. A Bloom Filter requires $k$ independent hash functions $h_{i}$ such that the range of each $h_{i}$ is the set of integers $\{1,\ldots,m\}$ [7].
Most mathematical treatments such as Mitzenmacher and Broder [7] make the convenient assumption that each $h_{i}$ maps each item in the universe to a random number uniformly over the (integer) range $[1,m]$. In the remainder of this text, we shall refer to this formulation of the Bloom Filter as the Classical Bloom Filter. This is to distinguish it from the Learned Bloom Filter, which our results are focused on.
@@
Figure 2: Example of a Learned Bloom filter with a Learned Oracle and a Backup Classical Bloom filter that only checks values that are with high probability negative in the Learned Oracle, to ensure a one-sided error bound (i.e only false positives and no false negatives).
@@ 1.1.2 Learned Bloom Filter
The Learned Bloom Filter is a novel data structure proposed by Kraska et al [1] in 2017. Kraska et al suggest using a pre-filter ahead of the Classical Bloom Filter, where the pre-filter is derived from a learning model. The learning model estimates the probability of an element $x$ being in the set $S$. This allows the use of a smaller (in terms of memory) Classical Bloom Filter compared to the case where a Classical Bloom Filter is used alone. A mathematical model and guarantees for the Learned Bloom Filter were first provided by Mitzenmacher [8] in 2018. While Kraska et al used a neural network for the pre-filter, Mitzenmacher generalizes the pre-filter to use any approach that estimates the probability of an element $x$ being in the set $S$. We use this generalized formulation of the Learned Bloom Filter in this work. We introduce the Learned Bloom Filter using a self-contained definition adapted from Mitzenmacher [8]. We extend this definition in Section 2.2.
@@
The Learned Bloom Filter provides better performance on the false positive rate while maintaining the guarantee of having no false negatives. We show an example of a Learned Bloom Filter adapted from Mitzenmacher et al [8] in figure 2.
@@ 1.2 Our Results
We introduce an adversarial model for the Learned Bloom Filter. Our model extends the adversarial model for the Classical Bloom Filter introduced by Naor Yogev [3]. Using our adversarial model, we introduce the notion of an adversarial resilient Learned Bloom Filter, as well as an adversarial reveal resilient Learned Bloom Filter. The precise definitions and the adversarial model we introduce are provided in Section 2.
@@ 1.2.1 Secure Learned Bloom Filter
We introduce the first provably secure constructions of the Learned Bloom Filter and show that using pseudorandom permutations, one can create provably secure constructions of the Learned Bloom Filter with at most $2\lambda$ extra bits of memory, where $\lambda$ is the security parameter.
@@
Let $B$ be an $(n,\epsilon)$-Sandwiched Learned Bloom Filter using $m$ bits of memory. If pseudo-random permutations exist, we show that
  1. 1.
    There exists a negligible function $\mathrm{negl}(\cdot)$ such that for security parameter $\lambda$ there exists an $(n,\epsilon+\mathrm{negl}(\lambda))$-adversarial resilient Learned Bloom Filter that uses $m^{\prime}=m+\lambda$ bits of memory
  2. 2.
    There exists a negligible function $\mathrm{negl}(\cdot)$ such that for security parameter $\lambda$ there exists an $(n,\epsilon+\mathrm{negl}(\lambda))$-adversarial reveal resilient Learned Bloom Filter that uses $m^{\prime}=m+2\lambda$ bits of memory
These results are formally proved in Theorem 3.7 and Theorem 3.10 of Section 3.
@@ 1.2.2 Utility of the Secure Learned Bloom Filter
$\alpha_{P}\text{FPR}(S^{\prime\prime},m_{A})+\alpha_{N}\text{FPR}(S^{\prime},m_{B})+(1-\alpha_{P}-\alpha_{N})\text{FPR}_{DB}(S,\mathcal{D},M)<\text{FPR}(S,m_{L}+m_{A}+m_{B}+\lambda)$
where $S^{\prime}\subset S$ is the set of elements in $S$ for which the Learned Oracle $L$ returns negative, $S^{\prime\prime}\subset S$ is the set of elements in $S$ for which the Learned Oracle $L$ returns positive, where the $FPR$ is a function for the false positive probability of a Secure Classical Bloom Filter, and $FPR_{DB}$ is the function for the false positive probability of a Secure Learned Bloom Filter.
8798These results are formally proven in Theorem 5.3 of Section 5. In this result, when we refer to a “Secure Learned Bloom Filter”, we refer in particular to the “Downtown Bodega Filter” construction we define in Section 3.3. We demonstrate multiple settings in a hybrid adversarial model where the Secure Learned Bloom Filter out-performs the Secure Classical Bloom Filter (Section 5.4). The advantage of the Secure Learned Bloom Filter compared to the Secure Classical Bloom Filter intricately depends on the number of adversarial queries and the performance of the Learned Oracle.
@@ 1.3 Motivation
The Bloom Filter and its variants have numerous applications in computing [9] [7]. We borrow discussion on Bloom Filter applications from the survey by Tarkoma et al [9]. The Bloom Filter may be implemented in kernel space in a Linux network driver for performant filtering of network packets. Loop detection in network protocols, and multicast forwarding engines may also utilize the Bloom Filter. Deep Packet Scanners and Packet Classifiers have also found the Bloom Filter helpful for improving efficiency. The Bloom Filter may be used to detect heavy flows in network traffic from the vantage point of a router. The Bloom Filter has also been used in the OPUS system [9] that stores a list of words that involve poor password choices encouraging users to select better passwords. The Bloom Filter has also found success in the detection of hash tampering in network attached disks. Google’s BigTable system uses the Bloom Filter to minimize disk reads. Apache Hadoop also uses the Bloom Filter as an optimization in the reduce stage of its map/reduce implementation. Other applications of the Bloom Filter include uses in the realms of peer-to-peer networking, and caching.
1.3.1 The Need for Security
A large number of the applications of the Bloom Filter involve critical infrastructure [10]. It is possible to forge false positives in a naively implemented Bloom Filter [10] allowing an adversary to make the Bloom Filter deviate from its behavior. Gerbet et al [10] show practical attacks on the Scrapy web-spider, the Bitly Dablooms spam filter, and the Squid web cache. Naor and Yogev [3] motivate the need for securing the Bloom Filter by considering a white-list of email addresses for the purposes of spam filtering. In their scenario, an adversary that can forge false positives may easily infiltrate the spam filter.
@@ 1.4.1 Classical Bloom Filter
Gerbet et al [10] suggest practical attacks on the Classical Bloom Filter and the use of universal hash functions and message authentication codes (MACs) to mitigate a subset of those attacks. Naor and Yogev [3] define an adversarial model for the Classical Bloom Filter and use it to prove that $(1)$ for computationally bounded adversaries, non-trivial adversary resilient Bloom filters exist if and only if one-way functions exist, and $(2)$ for computationally unbounded adversaries, there exists a Classical Bloom Filter that is secure against $t$ queries while using only $\mathcal{O}(n\log\frac{1}{\epsilon}+t)$ bits of memory. $n$ is the size of the set and $\epsilon$ is the desired error. We borrow their idea of using Pseudorandom Permutations for the Classical Bloom Filter and apply it to the Learned Bloom Filter.
Clayton et al [11] analyze not only the Classical Bloom Filter, but also the Counting Bloom Filter, and the Count-Min Sketch, in an adversarial setting. Clayton et al use a stronger adversarial model than Naor and Yogev [3], allowing an adversary to perform insertions and giving an adversary access to the internal state of the Classical Bloom Filter. Clayton et al propose the use of salts and keyed pseudo-random functions for securing the Classical Bloom Filter. They do not address Learned Probabilistic Data Structures including the Learned Bloom Filter. Both Naor and Yogev, and Clayton et al, perform their analysis in a game-based setting.
Filic et al [12] investigate the adversarial correctness and privacy of the Classical Bloom Filter and an insertion-only variant of the Cuckoo Filter. They use a stronger adversarial model than Naor and Yogev [3] allowing an adversary to insert entries into the Classical Bloom Filter and query for the internal state of the Classical Bloom Filter. Unlike our work, Filic et al [12] do not address the adversarial correctness of the Learned Bloom Filter or its variants. Filic et al [12] perform their analysis in a simulator-based setting.
@@ 1.4.2 Learned Bloom Filter
The authors are only aware of one prior work that addresses the Learned Bloom Filter in an adversarial setting, Reviriego et al [2]. They propose a practical attack on the Learned Bloom Filter. They suggest two possible mitigations for their proposed attack: swapping to a Classical Bloom Filter upon detection of the attack, or adding a second Backup Classical Bloom Filter. However, they do not provide any provable guarantees on the performance of the Learned Bloom Filter in the presence of adversaries. They leave the security of the Learned Bloom Filter as an open problem in their work.
@@ 2 Adversarial Model
We first describe the adversarial model of Naor and Yogev [3] for the Classical Bloom Filter and then use it as the basis for creating an adversarial model for the Learned Bloom Filter. We refer to the adversarial model defined by Naor and Yogev [3] as the classical adversarial model. Section 2.1 contains a treatment of the classical adversarial model. We also introduce a stronger adversary than the one described in Naor and Yogev’s [3] model that has access to the internal state of the Classical Bloom Filter. Section 2.2 introduces a definition of the Learned Bloom Filter adapted from [8] and discusses extensions to the classical adversarial model to make it work with the Learned Bloom Filter.
@@ 2.1 Classical Adversarial Model
We define a special class of Classical Bloom Filters which were coined “steady” Classical Bloom Filters by Naor and Yogev [3]. Steady Classical Bloom Filters do not change their internal representation $M^{S}_{r}$ after the setup algorithm $C_{r}$ has executed. In other words, only query algorithms of the type $Q_{s}$ are permitted in the steady setting and query algorithms of the type $Q_{u}$ are not permitted.
@@
Now, we create an extension to Naor and Yogev’s [3] model, introducing a stronger adversary that has access to the internal state of the Classical Bloom Filter. We construct our second adversarial challenge for the Classical Bloom Filter in the steady setting. This challenge is almost identical to the first challenge with the only difference being that the adversary is allowed access to the representation $M_{r}^{S}$. We define our second challenge more precisely in Challenge 2.
@@ 2.2 Learned Adversarial Model
Mitzenmacher [8] was the first to create a mathematical model for the Learned Bloom Filter, in 2018. Our model is heavily based on Mitzenmacher’s model but with some additional definitions to suit our adversarial setting. Consider a set of elements $\mathcal{K}\subset S$ and a set of elements $\mathcal{U}$ such that $\forall u\in\mathcal{U},u\notin S$. We form a training dataset $\mathcal{D}=\{(x_{i},y_{i}=1)|x_{i}\in\mathcal{K}\}\cup\{(x_{i},y_{i}=0)|x_{i}\in\mathcal{U}\}$.
@@
We train a Bloom Filter Learning Model, $l_{r}$, on $\mathcal{D}$. Let $l_{r}(x)$ be the probability estimate from the learning model that $x$ is an element in $S$. A value $\tau$ may be chosen as a threshold. When $l(x)\geq\tau$ then the Learned Bloom Filter considers $x$ to be an element of $S$. Otherwise, the Learned Bloom Filter passes $x$ on to the Backup Classical Bloom Filter. Figure 2 provides a helpful illustration. In many places in the text, we refer to operations involving $l_{r}$ and $\tau$ as operations of a Learned Oracle. This is to abstract away from the exact details of the Learning Model when it is unnecessary for the results we are trying to prove. We define this term precisely here, to avoid confusion.
Definition 2.10 (Learned Oracle).
Let a Learned Oracle $L$ be an oracle that, given a Bloom Filter Learning Model $l_{r}$ and a threshold $\tau$, returns positive if $l_{r}(x)\geq\tau$. $L$ returns negative otherwise.

### v3

@@ Abstract
Creating an adversary resilient construction of the Learned Bloom Filter with provable guarantees is an open problem. We define a strong adversarial model for the Learned Bloom Filter. Our adversarial model extends an existing adversarial model designed for the Classical (i.e not “Learned”) Bloom Filter by prior work and considers computationally bounded adversaries that run in probabilistic polynomial time (PPT). Using our model, we construct an adversary resilient variant of the Learned Bloom Filter called the Downtown Bodega Filter. We show that: if pseudo-random permutations exist, then an Adversary Resilient Learned Bloom Filter may be constructed with $2\lambda$ extra bits of memory and at most one extra pseudo-random permutation in the critical path. We construct a hybrid adversarial model for the case where a fraction of the query workload is chosen by an adversary. We show realistic scenarios where using the Downtown Bodega Filter gives better performance guarantees compared to alternative approaches in this hybrid model.
keywords:
Secret Key Cryptography, Adversarial Artificial Intelligence, Probabilistic Data Structures.
{textabstract}
Creating an adversary resilient construction of the Learned Bloom Filter with provable guarantees is an open problem. We define a strong adversarial model for the Learned Bloom Filter. Our adversarial model extends an existing adversarial model designed for the Classical (i.e not “Learned”) Bloom Filter by prior work and considers computationally bounded adversaries that run in probabilistic polynomial time (PPT). Using our model, we construct an adversary resilient variant of the Learned Bloom Filter called the Downtown Bodega Filter. We show that: if pseudo-random permutations exist, then an Adversary Resilient Learned Bloom Filter may be constructed with 2λ extra bits of memory and at most one extra pseudo-random permutation in the critical path. We construct a hybrid adversarial model for the case where a fraction of the query workload is chosen by an adversary. We show realistic scenarios where using the Downtown Bodega Filter gives better performance guarantees compared to alternative approaches in this hybrid model.
@@ 1 Introduction
The Bloom Filter is a probabilistic data structure that solves the Approximate Membership Query Problem. The data structure now known as the “Bloom” Filter was initially proposed as method $2$ in the section “Two Hash-Coding Methods with Allowable Errors” in a 1970 paper by Burton H. Bloom [Blo70][CRJ10]. The Bloom Filter has applications in databases, cryptography, computer networking, social networking [BGK+08], and network security [MB04]. The Learned Bloom Filter is a novel data structure first proposed by Kraska et al [KBC+18] in 2017. The Learned Bloom Filter can be thought of as a Bloom Filter working in collaboration with a Learning Model. There are currently no known provably secure constructions of the Learned Bloom Filter. Prior work [RAHDS21] has left the security of the Learned Bloom Filter as an open problem.
@@ 1.1.1 Classical Bloom Filter
A Bloom Filter for representing set $S$ with cardinality $n$ is a zero initialized array of $m$ bits. A Bloom Filter requires $k$ independent hash functions $h_{i}$ such that the range of each $h_{i}$ is the set of integers $\{1,\ldots,m\}$ [MB04].
Most mathematical treatments such as Mitzenmacher Broder [MB04] make the convenient assumption that each $h_{i}$ maps each item in the universe to a random number uniformly over the (integer) range $[1,m]$. In the remainder of this text, we shall refer to this formulation of the Bloom Filter as the Classical Bloom Filter. This is to distinguish it from the Learned Bloom Filter, which our results are focused on.
@@
Figure 2: Example of a Learned Bloom filter with a Learning Model and a Backup Classical Bloom filter that only checks values that are with high probability negative in the Bloom Filter Learning Model, to ensure a one-sided error bound (i.e only false positives and no false negatives).
@@ 1.1.2 Learned Bloom Filter
The Learned Bloom Filter is a novel data structure proposed by Kraska et al [KBC+18] in 2017. Kraska et al suggest using a pre-filter ahead of the Classical Bloom Filter, where the pre-filter is derived from a learning model. The learning model estimates the probability of an element $x$ being in the set $S$. This allows the use of a smaller (in terms of memory) Classical Bloom Filter compared to the case where a Classical Bloom Filter is used alone. A mathematical model and guarantees for the Learned Bloom Filter were first provided by Mitzenmacher [Mit18] in 2018. While Kraska et al used a neural network for the pre-filter, Mitzenmacher generalizes the pre-filter to use any approach that estimates the probability of an element $x$ being in the set $S$. We use this generalized formulation of the Learned Bloom Filter in this work. We introduce the Learned Bloom Filter using a self-contained definition adapted from Mitzenmacher [Mit18]. We extend this definition in Section 2.2.
@@
The Learned Bloom Filter provides better performance on the false positive rate while maintaining the guarantee of having no false negatives. We show an example of a Learned Bloom Filter adapted from Mitzenmacher et al [Mit18] in figure 2.
@@ 1.2 Our Results
We introduce an adversarial model for the Learned Bloom Filter. Our model extends the adversarial model for the Classical Bloom Filter introduced by Naor Yogev [NE19]. Using our adversarial model, we introduce the notion of an adversarial resilient Learned Bloom Filter, as well as an adversarial reveal resilient Learned Bloom Filter. The precise definitions and the adversarial model we introduce are provided in Section 2.
@@ 1.2.1 Secure Learned Bloom Filter
We introduce the first provably secure constructions of the Learned Bloom Filter and show that using pseudo-random permutations, one can create provably secure constructions of the Learned Bloom Filter with at most $2\lambda$ extra bits of memory, where $\lambda$ is the security parameter.
@@
Let $B$ be an $(n,\epsilon)$-Sandwiched Learned Bloom Filter using $m$ bits of memory. If pseudo-random permutations exist, we show that there exists a negligible function $\mathrm{negl}(\cdot)$ such that for security parameter $\lambda$ there exists an $(n,\epsilon+\mathrm{negl}(\lambda))$-adversarial reveal resilient Learned Bloom Filter that uses $m^{\prime}=m+2\lambda$ bits of memory
This result is formally proved in Theorem 3.7 of Section 3.
@@ 1.2.2 Utility of the Secure Learned Bloom Filter
$\begin{split}\alpha_{P}\text{FPR}(S^{\prime\prime},m_{A})+\alpha_{N}\text{FPR}(S^{\prime},m_{B})+(1-\alpha_{P}-\alpha_{N})\text{FPR}_{DB}(S,\mathcal{D},M)\\ <\text{FPR}(S,m_{L}+m_{A}+m_{B}+\lambda)\end{split}$ (1)
where $S^{\prime}\subset S$ is the set of elements in $S$ for which the Bloom Filter Learning Model $L$ returns negative, $S^{\prime\prime}\subset S$ is the set of elements in $S$ for which the Bloom Filter Learning Model $L$ returns positive, where the $FPR$ is a function for the false positive probability of a Secure Classical Bloom Filter, and $FPR_{DB}$ is the function for the false positive probability of a Secure Learned Bloom Filter.
These results are formally proven in Theorem 5.5 of Section 5. In this result, when we refer to a “Secure Learned Bloom Filter”, we refer in particular to the “Downtown Bodega Filter” construction we define in Section 3.2. We demonstrate multiple settings in a hybrid adversarial model where the Secure Learned Bloom Filter out-performs the Secure Classical Bloom Filter (Section 5.4). The advantage of the Secure Learned Bloom Filter compared to the Secure Classical Bloom Filter intricately depends on the number of adversarial queries and the performance of the Bloom Filter Learning Model.
@@ 1.3 Motivation
The Bloom Filter and its variants have numerous applications in computing [TRL12] [MB04]. We borrow discussion on Bloom Filter applications from the survey by Tarkoma et al [TRL12]. The Bloom Filter may be implemented in kernel space in a Linux network driver for performant filtering of network packets. Loop detection in network protocols, and multicast forwarding engines may also utilize the Bloom Filter. Deep Packet Scanners and Packet Classifiers have also found the Bloom Filter helpful for improving efficiency. The Bloom Filter may be used to detect heavy flows in network traffic from the vantage point of a router. The Bloom Filter has also been used in the OPUS system [TRL12] that stores a list of words that involve poor password choices encouraging users to select better passwords. The Bloom Filter has also found success in the detection of hash tampering in network attached disks. Google’s BigTable system uses the Bloom Filter to minimize disk reads. Apache Hadoop also uses the Bloom Filter as an optimization in the reduce stage of its map/reduce implementation. Other applications of the Bloom Filter include uses in the realms of peer-to-peer networking, and caching.
A large number of the applications of the Bloom Filter involve critical infrastructure [GKL15]. It is possible to forge false positives in a naively implemented Bloom Filter [GKL15] allowing an adversary to make the Bloom Filter deviate from its behavior. Gerbet et al [GKL15] show practical attacks on the Scrapy web-spider, the Bitly Dablooms spam filter, and the Squid web cache. Naor Yogev [NE19] motivate the need for securing the Bloom Filter by considering a white-list of email addresses for the purposes of spam filtering. In their scenario, an adversary that can forge false positives may easily infiltrate the spam filter.
@@ 1.4.1 Classical Bloom Filter
Gerbet et al [GKL15] suggest practical attacks on the Classical Bloom Filter and the use of universal hash functions and message authentication codes (MACs) to mitigate a subset of those attacks. Naor Yogev [NE19] define an adversarial model for the Classical Bloom Filter and use it to prove that $(1)$ for computationally bounded adversaries, non-trivial adversary resilient Bloom filters exist if and only if one-way functions exist, and $(2)$ for computationally unbounded adversaries, there exists a Classical Bloom Filter that is secure against $t$ queries while using only $\mathcal{O}(n\log\frac{1}{\epsilon}+t)$ bits of memory. $n$ is the size of the set and $\epsilon$ is the desired error. We borrow their idea of using Pseudorandom Permutations for the Classical Bloom Filter and apply it to the Learned Bloom Filter.
Clayton et al [CPS19] analyze not only the Classical Bloom Filter, but also the Counting Bloom Filter, and the Count-Min Sketch, in an adversarial setting. Clayton et al use a stronger adversarial model than Naor Yogev [NE19], allowing an adversary to perform insertions and giving an adversary access to the internal state of the Classical Bloom Filter. Clayton et al propose the use of salts and keyed pseudo-random functions for securing the Classical Bloom Filter. They do not address Learned Probabilistic Data Structures including the Learned Bloom Filter. Both Naor Yogev, and Clayton et al, perform their analysis in a game-based setting.
Filic et al [FPUV22] investigate the adversarial correctness and privacy of the Classical Bloom Filter and an insertion-only variant of the Cuckoo Filter. They use a stronger adversarial model than Naor Yogev [NE19] allowing an adversary to insert entries into the Classical Bloom Filter and query for the internal state of the Classical Bloom Filter. Unlike our work, Filic et al [FPUV22] do not address the adversarial correctness of the Learned Bloom Filter or its variants. Filic et al [FPUV22] perform their analysis in a simulator-based setting.
@@ 1.4.2 Learned Bloom Filter
The authors are only aware of one prior work that addresses the Learned Bloom Filter in an adversarial setting, Reviriego et al [RAHDS21]. They propose a practical attack on the Learned Bloom Filter. They suggest two possible mitigations for their proposed attack: swapping to a Classical Bloom Filter upon detection of the attack, or adding a second Backup Classical Bloom Filter. However, they do not provide any provable guarantees on the performance of the Learned Bloom Filter in the presence of adversaries. They leave the security of the Learned Bloom Filter as an open problem in their work.
@@ 2 Adversarial Model
We first describe the adversarial model of Naor Yogev [NE19] for the Classical Bloom Filter and then use it as the basis for creating an adversarial model for the Learned Bloom Filter. We refer to the adversarial model defined by Naor Yogev [NE19] as the classical adversarial model. Section 2.1 contains a treatment of the classical adversarial model. We also introduce a stronger adversary than the one described in Naor Yogev’s [NE19] model that has access to the internal state of the Classical Bloom Filter. Section 2.2 introduces a definition of the Learned Bloom Filter adapted from [Mit18] and discusses extensions to the classical adversarial model to make it work with the Learned Bloom Filter.
@@ 2.1 Classical Adversarial Model
We define a special class of Classical Bloom Filters which were coined “steady” Classical Bloom Filters by Naor Yogev [NE19]. Steady Classical Bloom Filters do not change their internal representation $M^{S}_{r}$ after the setup algorithm $C_{r}$ has executed. In other words, only query algorithms of the type $Q_{s}$ are permitted in the steady setting and query algorithms of the type $Q_{u}$ are not permitted.
@@
Now, we create an extension to Naor Yogev’s [NE19] model, introducing a stronger adversary that has access to the internal state of the Classical Bloom Filter. We construct our second adversarial challenge for the Classical Bloom Filter in the steady setting. This challenge is almost identical to the first challenge with the only difference being that the adversary is allowed access to the representation $M_{r}^{S}$. We define our second challenge more precisely in Challenge 2.
@@ 2.2 Learned Adversarial Model
Mitzenmacher [Mit18] was the first to create a mathematical model for the Learned Bloom Filter, in 2018. Our model is heavily based on Mitzenmacher’s model but with some additional definitions to suit our adversarial setting. Consider a set of elements $\mathcal{K}\subset S$ and a set of elements $\mathcal{U}$ such that $\forall u\in\mathcal{U},u\notin S$. We form a training dataset $\mathcal{D}=\{(x_{i},y_{i}=1)|x_{i}\in\mathcal{K}\}\cup\{(x_{i},y_{i}=0)|x_{i}\in\mathcal{U}\}$.
@@
We train a Bloom Filter Learning Model, $l_{r}$, on $\mathcal{D}$. Let $l_{r}(x)$ be the probability estimate from the learning model that $x$ is an element in $S$. A value $\tau$ may be chosen as a threshold. When $l(x)\geq\tau$ then the Learned Bloom Filter considers $x$ to be an element of $S$. Otherwise, the Learned Bloom Filter passes $x$ on to the Backup Classical Bloom Filter. Figure 2 provides a helpful illustration.